OpenStudy (anonymous):

prove (1+sinx)/tanx=cosx+cotx ??

OpenStudy (anonymous):

(1+sinx)/tanx = 1/tan(x) + sin(x)/tan(x) 1/tan(x) = cot(x) 1/tan(x) = cos(x)/sin(x); 1/tan(x) * sin(x) = cos(x)